# MR-EEGWaveNet: Multiresolutional EEGWaveNet for Seizure Detection from Long EEG Recordings

### v1

@@ 6.2 Effect of Window Length on Model Performance
Furthermore, we investigated our MR-EEGWaveNet with two different window length sizes and three different multiresolutional parameters, which are as follows: MR-EEGWaveNet-1 [5 s, 2.5 s], MR-EEGWaveNet-2 [10 s, 2 s], and MR-EEGWaveNet-3 [10 s, 5 s, 2 s]. The window length of model MR-EEGWaveNet-1 was $5$ s, where both MR-EEGWaveNet-2 and MR-EEGWaveNet-3 were $10$ s long. For the Siena dataset (see Table 4), all metrics except recall improved with longer window lengths, whereas recall decreased by approximately 2% in the MR-EEGWaveNet-2 and MR-EEGWaveNet-3 models compared to MR-EEGWaveNet-1. For the Juntendo dataset (see Table 5), all improved metrics remain almost the same with longer window lengths.
@@ 6.4 Contribution of the Post-Classification Processing
[…]
@@ 6.5 Feature Visualization
Figure 3: t-SNE visualization of features calculated from EEG recordings from the EEGWaveNet and the MR-EEGWaveNet. (Dataset: Juntendo, Patient ID: $12$)
Furthermore, we used the t-distributed stochastic neighbor embedding (t-SNE) technique to visualize the distribution of features in two dimensions, as shown in Figure 3. The features were extracted from the final layer of both the EEGWaveNet and the MR-EEGWaveNet.
@@ 7.3 Performance Improvement with MR-EEGWaveNet
The patient-wise performance of the EEGWaveNet-3 and MR-EEGWaveNet-2 on the Siena dataset with post-classification processing is shown in Tables A1 and A2, respectively. Furthermore, Tables B1 and B2 represent the results on the Juntendo dataset for the EEGWaveNet-3 and MR-EEGWaveNet-2 methods, respectively. Notably, the MR-EEGWaveNet demonstrates consistent performance across multiple datasets, further supporting its scalability and stability. However, more studies with larger datasets are necessary to confirm this finding.
@@ 7.6 Comparison with Recent Studies
The Siena dataset utilized in this study is a relatively new dataset available publicly in the field. To the best of our knowledge, among the limited number of recent studies utilizing the Siena data set, the work by Sigsgaard et al. [25], which adopts a leave-one-subject-out evaluation protocol, represents the most comparable baseline for evaluating our approach. Although the method proposed in [25] is segment-based, the performance evaluation reported in their study is event-based. Table 8 presents a performance comparison between MR-EEGWaveNet-2 and the event-based approach proposed by Sigsgaard et al. In terms of performance based on seizures, our proposed method outperforms the approach of Sigsgaard et al. [25] by approximately $13\%$ in the detection of seizures, with a seizure event detection rate of $\sim 94\%$. However, in terms of specificity, it falls short by approximately $4\%$. Therefore, our proposed method effectively detects seizures in both segment-wise and event-based evaluation schemes, while maintaining a moderate level of specificity and a false positive rate. Furthermore, note that models designed explicitly for seizure event detection generally exhibit higher specificity compared to segment-based approaches. However, the performance reported largely depends on how seizure detection is defined and evaluated. In this study, a seizure event is considered detected if the model labels at least one EEG segment within the seizure period as a seizure, regardless of whether all segments of the event are labeled. For example, if an annotated seizure occurs between 100 and 150 seconds and the EEG is divided into 10-second segments, the event is considered detected if the model classifies any overlapping segment (e.g., 100–110 s or 110–120 s) as a seizure, even if not all segments within the 100–150 s period are detected.
@@ 7.8 Limitations and Future Directions
A notable limitation of our model is its reduced performance for certain subjects. Further improvements are needed to enhance the model’s ability to distinguish seizures from artifacts, such as loss of contact with the electrode, eye movements (EOG), or muscle activity. In the future, the model architecture should include context-based feature extraction to better capture temporal and spatial patterns. Furthermore, improving the model’s overall reliability and performance will require the application of advanced artifact removal and post-classification processing techniques.
@@ 8 Conclusion
This study presents a novel end-to-end model to detect epileptic seizures from long EEG recordings. The length of the EEG segment plays a significant role in the model’s performance, especially in distinguishing between seizure, noise, and artifact. The experimental results showed that the MR-EEGWaveNet with multiresolutional features performed better than the existing end-to-end models in classifying seizure and nonseizure EEG signals. Additionally, we introduced an anomaly score-based post-classification processing technique to further improve performance and reduce false positive rates. Furthermore, we demonstrated the performance of the MR-EEGWaveNet with different parameter settings. Despite lower performance in certain subjects, the improvement achieved with MR-EEGWaveNet indicates its potential for further development. However, enhancing MR-EEGWaveNet is essential to improve its ability to effectively distinguish between artifacts and seizure events. Although the MR-EEGWaveNet was tested on two datasets and demonstrated consistent performance, further analysis and investigation with larger datasets are required to apply it in real-world applications.
Appendix A Siena Dataset: Patient-Wise Results
A.1 EEGWaveNet Model
Table A1: EEGWaveNet-3, Patient-wise performance (Window length: $10$ s), with post-classification processing on Siena dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
@@ A.1 EEGWaveNet Model
A.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table A2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Siena dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
@@ A.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Appendix B Juntendo Dataset: Patient-wise Results
B.1 EEGWaveNet Model
Table B1: EEGWaveNet-3, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
@@ B.1 EEGWaveNet Model
B.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table B2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.

### v2

@@ 6.2 Effect of Window Length on Model Performance
Furthermore, we investigated our MR-EEGWaveNet with two different window length sizes and three different multiresolutional parameters, which are as follows: MR-EEGWaveNet-1 [5 s, 2.5 s], MR-EEGWaveNet-2 [10 s, 2 s], and MR-EEGWaveNet-3 [10 s, 5 s, 2 s]. The window length of model MR-EEGWaveNet-1 was $5$ s, where both MR-EEGWaveNet-2 and MR-EEGWaveNet-3 were $10$ s long. For the Siena dataset (see Table 4), all metrics except recall improved with longer window lengths, while recall decreased by approximately 2% in the MR-EEGWaveNet-2 and MR-EEGWaveNet-3 models compared to MR-EEGWaveNet-1. For the Juntendo dataset (see Table 5), similar improvements in performance were observed with longer window lengths.
@@ 6.4 Contribution of the Post-Classification Processing
Figure 4: Outcome of the MR-EEGWaveNet-2 model on a 1-hour EEG recording after post-classification processing.
@@
In Table 7, we present a similar comparative study between EEGWaveNet and MR-EEGWaveNet in the Juntendo dataset. The experimental results demonstrated performance improvements across all metrics (except recall, which remains almost the same), similar to the Siena dataset. The post-classification processing improved the specificity of the proposed model by approximately $3\%$, and by $7-8\%$ for the EEGWaveNet models (although recall dropped slightly by about $1-2\%$). Those performance improvements indicate the effectiveness of the post-classification processing technique with the EEGWaveNet and the proposed models.
@@ 6.5 Feature Visualization
Figure 5: t-SNE visualization of features extracted from EEG recordings using EEGWaveNet and MR-EEGWaveNet (Dataset: Juntendo; Patient ID: Pt-$12$).
Furthermore, we used the t-distributed stochastic neighbor embedding (t-SNE) technique to visualize the distribution of features in two dimensions, as shown in Figure 5. The features were extracted from the final layer before the prediction module in both the EEGWaveNet and the MR-EEGWaveNet models.
6.6 Ablation Study with MR-EEGWaveNet Variants
Table 8 presents the performance of two MR-EEGWaveNet variants and the baseline MR-EEGWaveNet-2 model on Siena dataset. Removing the feature extraction module (MR-EEGWaveNet-Abl1) significantly reduces performance, with a precision of $21.53\%$ and F1 score of $0.259$. MR-EEGWaveNet-Abl2, which excludes the $10$-second input stream, improves recall ($83.10\%$) but yields lower F1 ($0.288$) and AUC ($0.9364$) than the baseline. The baseline MR-EEGWaveNet-2 achieves the best performance across all metrics, including the highest precision ($26.74\%$), specificity ($93.86\%$) F1 score ($0.319$), and AUC ($0.9421$). The patient-wise performance details of the MR-EEGWaveNet-Abl1 and MR-EEGWaveNet-Abl2 models are reported in Appendix B, in Tables B3 and B4, respectively.
Table 8: Performance evaluation of MR-EEGWaveNet and ablated variants on the Siena dataset (best performances are highlighted in bold).
Models [Segments] Pre. Rec. Spe. F1 AUC
MR-EEGWaveNet-Abl1 [10 s] 21.53 81.23 91.75 0.259 0.9379
MR-EEGWaveNet-Abl2 [10 s] 23.17 83.10 93.04 0.288 0.9364
MR-EEGWaveNet-2 [10 s, 2 s] 26.74 82.39 93.86 0.319 0.9421
Table 9: Comparison of CBraMod (with LightGBM) and MR-EEGWaveNet performance on the Siena dataset.
Models [Segments] Pre. Rec. Spe. F1 AUC
CBraMod + LightGBM [10 s] 8.74 52.05 92.80 0.120 0.8756
MR-EEGWaveNet-2 [10 s, 2 s] 26.74 82.39 93.86 0.319 0.9421
6.7 Evaluation Against Pretrained Foundation Model
Table 9 presents a performance comparison between the proposed MR-EEGWaveNet-2 model and the CBraMod model combined with LightGBM. The MR-EEGWaveNet-2 consistently outperforms the CBraMod-based approach across all evaluation metrics. Specifically, it achieves substantially higher precision ($26.74\%$ vs. $8.74\%$), recall ($82.39\%$ vs. $52.05\%$), F1 score ($0.319$ vs. $0.120$), and AUC ($0.9421$ vs. $0.8756$). The patient-wise performance of CBraMod combined with the LightGBM classifier is reported in Appendix B, Table B5.
@@ 7.3 Performance Improvement with MR-EEGWaveNet
The patient-wise performance of the EEGWaveNet-3 and MR-EEGWaveNet-2 on the Siena dataset with post-classification processing is shown in Tables B1 and B2, respectively. Furthermore, Tables C1 and C2 represent the results on the Juntendo dataset for the EEGWaveNet-3 and MR-EEGWaveNet-2 methods, respectively. Notably, the MR-EEGWaveNet demonstrates consistent performance across multiple datasets, further supporting its scalability and stability. However, more studies with larger datasets are necessary to confirm this finding.
@@
7.4 Ablation Study on MR-EEGWaveNet Model Components
MR-EEGWaveNet-Abl1 excludes the feature extraction module, which significantly reduces the model’s ability to learn rich temporal and spatial representations. This simplification leads to the lowest performance across all metrics (see Table 8)–indicating that the feature extractor plays a critical role in distinguishing seizure segments.
MR-EEGWaveNet-Abl2 removes the $10$-second input stream, retaining only the $2$-second sub-segment pipeline. This variant performs better than MR-EEGWaveNet-Abl1 in all metrics and achieves the highest recall ($83.10\%$), suggesting that the $2$-second sub-segments alone still provide useful information. However, the drop in other metrics, specifically AUC and specificity compared to the baseline, indicates that the longer context ($10$-second input) contributes to overall decision stability and discriminative power with fewer false positives.
The baseline model, MR-EEGWaveNet-2, which integrates both the feature extraction module and the multi-scale input pipeline (10 s and 2 s), achieves the best performance across all metrics. These improvements indicate that both components are complementary and essential to the model’s effectiveness.
@@ 7.7 Comparison with Recent Studies
As shown in Table 9, the comparison with the CBraMod (with LightGBM) foundation model highlights the strengths of the proposed MR-EEGWaveNet-2 architecture in the context of seizure detection from long-term EEG recordings. Despite CBraMod’s large-scale pre-training and use of powerful foundation model representations, its performance falls significantly short of MR-EEGWaveNet-2 when applied to the target dataset via feature extraction and LightGBM classification. Notably, MR-EEGWaveNet-2 achieved much higher precision ($26.74\%$ vs. $8.74\%$) and recall ($82.39\%$ vs. $52.05\%$), indicating its effectiveness in correctly identifying seizure events while minimizing false positives. Moreover, improvements in F1 score ($0.319$ vs. $0.120$) and AUC ($0.9421$ vs. $0.8756$) confirm the model’s robustness against class imbalance and variability in EEG recordings. These findings indicate that end-to-end task-specific models like MR-EEGWaveNet outperform general pretrained models, especially in capturing domain-specific temporal and contextual features.
The Siena dataset utilized in this study is a relatively new dataset available publicly in this field. To the best of our knowledge, among the limited number of recent studies utilizing the Siena data set, the work by Sigsgaard et al. [25], which adopts LOSO evaluation protocol, represents the most comparable baseline for evaluating our approach. Although the method proposed in [25] is segment-based, the performance evaluation reported in their study is event-based. Table 10 presents a performance comparison between MR-EEGWaveNet-2 and the event-based approach proposed by Sigsgaard et al. In terms of performance based on seizures, our proposed method outperforms the approach of Sigsgaard et al. [25] by approximately $13\%$ in the detection of seizures, with a seizure event detection rate of $\sim 94\%$. However, in terms of specificity, it falls short by approximately $4\%$. Therefore, our proposed method effectively detects seizures in both segment-wise and event-based evaluation schemes, while maintaining a moderate level of specificity and a false positive rate. Furthermore, note that models designed explicitly for seizure event detection generally exhibit higher specificity compared to segment-based approaches. However, the performance reported largely depends on how seizure detection is defined and evaluated. In this study, a seizure event is considered detected if the model labels at least one EEG segment within the seizure period as a seizure, regardless of whether other segments of the seizure event are labeled. For example, if an annotated seizure occurs between 100 and 150 seconds and the EEG is divided into $10$-second segments, the event is considered detected if the model classifies any overlapping segment (e.g., 100–110 s or 110–120 s) as a seizure, even if not all segments within the 100–150 s period are detected.
@@ 7.9 Limitations and Future Directions
A notable limitation of our model is its reduced performance for certain subjects. Further improvements are needed to enhance the model’s ability to distinguish seizures from artifacts, such as loss of contact with the electrode, eye movements, or muscle activity. In the future, the model architecture should include context-based feature extraction to better capture temporal and spatial patterns. Furthermore, improving the model’s overall reliability and performance will require the application of advanced artifact removal and post-classification processing techniques.
@@ 8 Conclusion
This study presents a novel end-to-end model to detect epileptic seizures from long EEG recordings. The length of the EEG segment plays a significant role in the model’s performance, especially in distinguishing between seizure, noise, and artifact. The experimental results showed that the MR-EEGWaveNet with multiresolutional features outperforms existing end-to-end models in classifying seizure and nonseizure EEG signals. Additionally, we introduced an anomaly score-based post-classification processing technique to improve performance and reduce false positive rates. We also evaluated the performance of MR-EEGWaveNet under different parameter settings and with ablated versions of its architecture. Despite lower performance in certain subjects, the improvement achieved with MR-EEGWaveNet indicates its potential for continued development. However, enhancing MR-EEGWaveNet is necessary to better distinguish between artifacts and seizure events. Although the MR-EEGWaveNet was tested on two datasets and demonstrated consistent performance, further investigation with larger datasets is required to apply it in real-world applications.
Code Availability
The code used in this study is available at https://github.com/ttlabtuat/MR-EEGWaveNet.
Appendix A Ablated Variants of MR-EEGWaveNet
A.1 Architecture of MR-EEGWaveNet Ablation Model 1 (MR-EEGWaveNet-Abl1)
Figure A1: MR-EEGWaveNet-Abl1 architecture
A.2 Architecture of MR-EEGWaveNet Ablation Model 2 (MR-EEGWaveNet-Abl2)
Figure A2: MR-EEGWaveNet-Abl2 architecture
Appendix B Siena Dataset: Patient-Wise Results
B.1 EEGWaveNet Model
Table B1: EEGWaveNet-3, Patient-wise performance (Window: $10$ s), with post-classification processing on Siena dataset.
@@ B.1 EEGWaveNet Model
Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score; The performance for each metric is reported as the average of three runs per patient.
B.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table B2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Siena dataset.
@@ B.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
B.3 MR-EEGWaveNet Ablation Model 1 (MR-EEGWaveNet-Abl1)
Table B3: MR-EEGWaveNet-Abl1, Patient-wise performance (Window: $10$ s), without post-classification processing on Siena dataset.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
PN00 77.27 98.72 98.94 0.850 0.999 1.000
PN01 7.80 100.0 96.43 0.144 1.000 1.000
PN03 2.53 92.42 90.06 0.049 0.970 1.000
PN05 4.35 87.96 88.96 0.082 0.951 1.000
PN06 28.67 80.29 97.71 0.358 0.986 1.000
PN07 1.58 86.67 90.57 0.031 0.952 1.000
PN09 86.06 83.65 99.92 0.847 0.998 1.000
PN10 10.06 46.20 98.15 0.161 0.891 0.482
PN11 3.41 63.89 85.62 0.063 0.777 1.000
PN12 3.09 90.56 60.13 0.060 0.904 1.000
PN13 10.66 97.22 92.92 0.191 0.977 1.000
PN14 6.60 63.55 98.19 0.119 0.915 0.667
PN16 15.70 61.23 89.32 0.216 0.882 1.000
PN17 43.62 84.80 97.60 0.465 0.930 1.000
Mean 21.53 81.23 91.75 0.259 0.937 0.939
B.4 MR-EEGWaveNet Ablation Model 2 (MR-EEGWaveNet-Abl2)
Table B4: MR-EEGWaveNet-Abl2, Patient-wise performance (Window: $10$ s), without post-classification processing on Siena dataset.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
PN00 72.49 100.0 99.13 0.840 0.999 1.000
PN01 16.69 100.0 96.26 0.269 1.000 1.000
PN03 2.23 91.17 88.60 0.044 0.945 1.000
PN05 3.25 96.30 87.40 0.063 0.946 1.000
PN06 34.19 58.75 99.37 0.409 0.971 0.933
PN07 1.13 93.33 85.85 0.022 0.927 1.000
PN09 47.95 81.11 97.89 0.520 0.988 1.000
PN10 13.08 45.08 98.64 0.193 0.799 0.593
PN11 4.27 100.0 83.93 0.082 0.934 1.000
PN12 6.19 87.96 80.51 0.115 0.921 1.000
PN13 23.89 85.04 96.61 0.346 0.975 1.000
PN14 5.55 61.17 97.88 0.101 0.899 0.667
PN16 9.02 71.21 90.69 0.160 0.873 1.000
PN17 84.47 92.28 99.87 0.881 0.933 1.000
Mean 23.17 83.10 93.04 0.288 0.936 0.942
Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score; The performance for each metric is reported as the average of three runs per patient.
B.5 CBraMod: Criss-Cross Brain Foundation Model
Table B5: CBraMod+LightGBM, Patient-wise performance (Window: $10$ s), without post-classification processing on Siena dataset.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
PN00 33.29 35.02 98.45 0.339 0.897 0.867
PN01 6.16 48.48 98.41 0.109 0.949 1.000
PN03 1.90 96.97 87.27 0.037 0.980 1.000
PN05 0.45 22.22 79.74 0.009 0.631 0.556
PN06 3.22 12.70 98.06 0.051 0.808 0.400
PN07 0.97 100.0 80.51 0.019 0.923 1.000
PN09 5.50 20.16 97.92 0.086 0.850 0.556
PN10 24.48 30.21 99.64 0.270 0.888 0.296
PN11 5.18 93.33 90.06 0.098 0.935 1.000
PN12 6.99 57.55 91.72 0.125 0.872 1.000
PN13 2.78 17.10 95.69 0.048 0.828 0.667
PN14 3.79 53.85 97.57 0.071 0.835 0.500
PN16 7.26 89.39 85.47 0.134 0.951 1.000
PN17 20.32 51.71 98.63 0.289 0.913 1.000
Mean 8.74 52.05 92.80 0.120 0.875 0.774
Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score; The performance for each metric is reported as the average of three runs per patient.
Appendix C Juntendo Dataset: Patient-wise Results
C.1 EEGWaveNet Model
Table C1: EEGWaveNet-3, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset.
@@ C.1 EEGWaveNet Model
Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score; The performance for each metric is reported as the average of three runs per patient.
C.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table C2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset.
@@ C.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score; The performance for each metric is reported as the average of three runs per patient.
Appendix D LightGBM Classifier
Table D1: LightGBM parameters for training
Parameter Value
boosting_type gbdt
class_weight balanced
colsample_bytree 1.0
importance_type split
learning_rate 0.1
max_depth -1
min_child_samples 20
min_child_weight 0.001
min_split_gain 0.0
n_estimators 100
n_jobs None
num_leaves 31
objective None
random_state 42
reg_alpha 0.0
reg_lambda 0.0
subsample 1.0
subsample_for_bin 200000
subsample_freq 0